The length of a rectangle is 3 feet more than the width. If the perimeter is 194 feet, what are the length and the width?
Answer by Cromlix(4381)

Hi there,
Make the width = x
Length = x + 3
Perimeter = 2 x Length + 2 x Width
194 = 2(x + 3) + 2(x)
194 = 2x + 6 + 2x
Collect like terms:
2x + 2x = 194 - 6
4x = 188
x = 47
Width = 47 feet
Length = 50 feet
Hope this helps :-)